how do u solve this problem 1-5xsquared=-3x
Answer by mukhopadhyay(490)

1-5x^2=-3x
=>5x^2-3x-1=0
=>x=[3+sqrt(9+20)]/10] or x=[3-sqrt(9+20)]/10]
=>x=3/10 + sqrt(29)/10 or x=3/10 - sqrt(29)/10